A crate has a mass of 42 kg is placed on an 50° incline surface. The coefficient of friction between the crate and the surface is 0.85. If the crate is pushed down the incline at a rate of 2.25 m/s2, determine:

a) the normal force 

 

b) the frictional force 


c) the pulling force P 



1
Expert's answer
2022-05-22T14:38:25-0400

a)

"N=mg\cos 50°=270~N,"

b)

"F=\mu N=230~N,"

c)

"P=m(g+a)(\sin 50°-\mu \cos 50°)=95~N."